solve for each formula for the indicated variable
F=mv^2/r for r

Answer by xcentaur(357)

F=mv^2/r
Taking r to the other side,
F.r=mv^2
r=mv^2/F


Sidenote: such basic formula rendition really shouldn't be asked here.But since I wouldn't like you to find things hard later,I made an exception solving this.


Hope this helps,
best of luck.